OpenStudy (anonymous):

Simplify: (x - 8)2 x2 + 64 x2 - 16x + 64 x2 - 64x + 16 x2 - 16 Simplify: (5x - 10)2 10x2 - 20 25x2 - 100 25x2 - 100x + 20 25x2 - 100x + 100

OpenStudy (anonymous):

\[\large (a -b)^2 = a^2 + b^2 - 2ab\] Here a = x, b = 8.. Put and find the value..

OpenStudy (lgbasallote):

or if you hate memorizing formulas... \[\huge (x-8)^2 \implies (x-8)(x-8)\] you can use FOIL here

OpenStudy (anonymous):

x2 - 16 for the 1st one

OpenStudy (anonymous):

Don't tell answer directly and if you want to then tell then right answer @muhammad9t5

OpenStudy (anonymous):

ok sir.

OpenStudy (anonymous):

@cutenerd will you try atleast once..

OpenStudy (jiteshmeghwal9):

\[\LARGE{\color{red}{(x-8)(x-8)}}\]

OpenStudy (jiteshmeghwal9):

\[\LARGE{\color{red}{x(x-8)-8(x-8)}}\]

OpenStudy (lgbasallote):

these replies torture the poor epileptics :C

OpenStudy (jiteshmeghwal9):

r u saying to me @lgbasallote ????

OpenStudy (lgbasallote):

lol. you guys get conscious easily =))))